Foreverangel:

I really need help please I can't get this and it is due soon. Perform the indicated operation.

Foreverangel:

1 attachment
Foreverangel:

1 attachment
ScaryPerson:

Az are you gonna help or can i?

AZ:

(g o f)(x) means you plug in f(x) into g(x) (g o f)(x) = g(f(x)) let me show you an example let's say we have f(x) = x + 1 g(x) = 2x if we're trying to find (g o f)(x) that means we take whatever f(x) is equal to (which is x+1 in this example) and we plug it into g(x) pay attention to the colors \( (g ~o~ f)(x) = g(\color{red}{f(x)})\) \( f(x) = \color{red}{x+1}\) \(g(x) = 2x\) \( g(\color{red}{f(x)}) = g(x+1)\) and we know that \(g(\color{green}{x}) = 2\color{green}{x}\) and we see that we have x+1 inside the parenthesis so we get \( g\color{green}{(x+1)} = 2\color{green}{(x+1})\) and simplifying that \( g(x+1) = 2x + 2\) so your final answer for this example would be (g o f)(x) = g(f(x)) = 2x + 2

AZ:

Remember, that was just an example but it's the same concept for your question (g o f)(x) = g(f(x)) what is f(x) equal to? and then replace that with 'x' in the g(x) equation and simplify

Foreverangel:

This is kind of how my teacher explained it I honestly just am not getting it sorry

Foreverangel:

I am trying to but I cant get it

AZ:

let's look at your question g(x) = x^2 + 5 f(x) = 4x - 3 what is f(x) = ??? hint: it's given to you in your question, you just need to restate it

Foreverangel:

4x-3?

AZ:

@foreverangel wrote:
4x-3?
Yes. Now we plug that into g(x) g(x) = x^2 + 5 we said f(x) is 4x-3 that means \( g(\color{red}{f(x)})\) is equal to \(g(\color{red}{4x-3})\) so we have \(g(\color{green}{x}) = \color{green}{x}^2 + 5\) replace the green 'x' with the red 4x-3 in both places and simplify

AZ:

hint: (a-b)^2 = a^2 - 2ab + b^2